OpenStudy (anonymous):

Center of a circle with this equation: y^2-6x-8y+20=-x^2

OpenStudy (mertsj):

Get the variables on the left and the constant term on the right. complete the square twice and write the equation in the form: \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

I did but I end up with what I believe is the wrong answer. Okay, so obviously it would be (x-4)^2+(y-3)^3=-20

OpenStudy (anonymous):

I am confused as to what to do with the neg.

OpenStudy (mertsj):

What negative?

OpenStudy (anonymous):

The neg in front of the twenty

OpenStudy (callisto):

y^2-6x-8y+20=-x^2 x^2 +y^2-6x-8y+20= 0 centre = (-D/2 , -E/2) in the form x^2 +y^2+Dx+Ey+F= 0 Therefore centre = ( -(-6/2) , -(-8/2) ) =...... Can you do it now?

OpenStudy (mertsj):

Nothing. It will be gone because you will be adding numbers to the right and the left sides.

OpenStudy (mertsj):

Ok. callisto wants to help you now. Good bye and good luck.

OpenStudy (anonymous):

center is therefore.. (-3,-4) right?

OpenStudy (callisto):

No... the sign is not correct

OpenStudy (anonymous):

sooo it's (3,4).

OpenStudy (anonymous):

So basically it's still the same center as it would be without the negatives...?

OpenStudy (callisto):

Yes

OpenStudy (anonymous):

Okay. Also, what the radius would be sqrtroot of 20.. right?

OpenStudy (anonymous):

Sorry, but just for conformation.. It doesn't matter if there is a negative for the 'r', the center would still be the same.

OpenStudy (callisto):

Nope.. \[r = \sqrt { (D/2)^2 + (E/2)^2 -F}\]

OpenStudy (callisto):

D=-6, E=-8 , F=20 in your case. Put the numbers into it and solve, what would you get for r?

OpenStudy (anonymous):

squareroot of 5...

OpenStudy (callisto):

It should be that ...

OpenStudy (anonymous):

That.. doesn't make sense though. The radius is typically sqroot of \[(x-h)^2 + (y-k)^2=r^2\]

OpenStudy (callisto):

So, let's do the completing square once y^2-6x-8y+20=-x^2 x^2 + y^2-6x-8y+20= 0 x^2 -6x +y^2-8y+20 =0 (x^2 -6x +9 -9) +(y^2-8y +16 -16)+20 =0 (x^2 -6x +9) +(y^2-8y +16) -9 -16 +20 =0 (x-3)^2 + (y-4)^2 - 5 =0 (x-3)^2 + (y-4)^2 = 5 (x-3)^2 + (y-4)^2 =[ sqrt (5) ]^2 r = sqrt 5 It DOES make sense

OpenStudy (callisto):

Here is the proof...